OpenStudy (anonymous):

For what values of x if any will the matrix A=[3 0 0; 0 x 2; 0 2 x] have at least one repeated eigenvalue?

OpenStudy (anonymous):

does it have something to do with symmetry?

OpenStudy (anonymous):

intuition tells me 2 or -2....but im working it out to see if thats the case

myininaya (myininaya):

| 3-c 0 0| |0 x-c 2| |0 2 x-c| =(3-c)[(x-c)^2-4] 3-c=0=>c=3 (x-c)^2=4 => x-c=2 and x-c=-2=> x=2+c and x=-2+c ....

myininaya (myininaya):

x=2+3=5 or x=-2+3=1

OpenStudy (anonymous):

Why do you have c's in your work?

OpenStudy (anonymous):

those are his eigenvalues (lambda)

myininaya (myininaya):

c is the eigenvalue

myininaya (myininaya):

her*

OpenStudy (anonymous):

sry sry, those are her eigenvalues (lambda)

myininaya (myininaya):

someone might want to check my x's is that right?

OpenStudy (anonymous):

Oh haha okay but I got (c-3)(c^2-2cx+x^2-4)

OpenStudy (anonymous):

im on it, one sec

myininaya (myininaya):

i think it is right

OpenStudy (anonymous):

its it :)

OpenStudy (anonymous):

just finished checking it, i wish i knew a fast way to do problems like these >.>

OpenStudy (anonymous):

one question though once you get the characteristic polynomial how'd you know what values of x works?

myininaya (myininaya):

first i found c

myininaya (myininaya):

or lambda whatever you want to name your eingenvalues

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

which would have a x value in it right

myininaya (myininaya):

then i used c to find the x's the other other factor of the characteristic polynomial was (x-c)^2-4 i set this=0 x-c=2 x-c=-2 but c=3 so one x is 5 another possible x is x=1

myininaya (myininaya):

det(A-cI)=0

OpenStudy (anonymous):

Oh okay make sense now thanks for all the help

myininaya (myininaya):

nice problem :)

OpenStudy (anonymous):

very interesting, hadnt seen one like it before.